# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 3 Impossibility Results
In the first result we prove that two agents can not solve PerpExploration-BBH in the strongest communication model of whiteboard.
@@
A set of two synchronous agents in a ring $R$ of size $n$ cannot solve PerpExploration-BBH, even in presence of whiteboard if number of possible consecutive black hole positions is greater or equals to 3.
Proof 3.2.
Let $v_{1},v_{2}$ and $v_{3}$ be three possible consecutive black hole positions in a ring $R$. Let two agents $a_{1}$ and $a_{2}$ are sufficient to solve the PerpExploration-BBH problem on $R$. Thus there exists an algorithm $\mathcal{A}$ such that $a_{1}$ and $a_{2}$ can solve PerpExploration-BBH by executing $\mathcal{A}$. Without loss of generality let $a_{1}$ explores $v_{2}$ and it moves to $v_{2}$ for the first time at round $t$ from vertex $v_{1}$. Then $a_{1}$ must have been at the vertex $v_{1}$ at round $t-1$. Let us take two copies $R_{1}$ and $R_{2}$ of the same ring $R$. In $R_{1}$, $v_{1}$ is the black hole and in $R_{2}$, $v_{2}$ is the black hole. Let the adversary in $R_{1}$ destroys $a_{1}$ at round $t-1$ and in $R_{2}$ destroys $a_{1}$ in round $t$. We claim that at round $t$ and $t-1$, $a_{2}$ can not be on either of $v_{1}$ or $v_{2}$. Note that at $t$-th round $a_{2}$ can not be on $v_{2}$ otherwise, both $a_{1}$ and $a_{2}$ gets destroyed at the same round and no other agent is left to explore further. Similarly at round $t-1$ $a_{2}$ can not be on $v_{1}$. Now consider the case when at $t-$ th round $a_{2}$ is at $v_{1}$. Since at round $t-1$, $a_{2}$ can not be on $v_{2}$ there can be two cases either at round $t-1$, $a_{2}$ was in $u_{0}$ ($u_{0}$ is another adjacent node of $v_{1}$ except $v_{2}$) or was in $v_{1}$ itself. We first argue that at round $t-1$, $a_{2}$ can not be at $u_{0}$. Otherwise, since the whiteboard content at nodes except at $v_{1}$ and $v_{2}$ are same in $R_{1}$ and $R_{2}$ at round $t$ (due to same execution by $a_{1}$ and $a_{2}$ upto round $t-1$ and at round $t$ for $a_{2}$ in both $R_{1}$ and $R_{2}$). Now since at round $t$ all previous datas on whiteboard except $v_{1}$ and $v_{2}$ does not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$, the problem now can be thought of solving PerpExploration-BBH with one agent where number of possible consecutive black hole position is atleast two. Now this is impossible to solve. So, at $t-1$-th round $a_{2}$ can not be on $u_{0}$ We now only have to prove that $a_{2}$ can not be at $v_{2}$ during round $t-1$. Let in $R_{2}$, adversary activates the black hole at $v_{2}$ at round $t-1$ which destroys $a_{2}$ at round $t-1$. Now since the effect of this destruction of agent $a_{2}$ does not effect the inputs of $a_{1}$ at round $t-1$ on $v_{1}$, it move to $v_{2}$ at round $t$ where the black hole is activated again by adversary destroying $a_{1}$ too. Thus $a_{2}$ must remain outside of $v_{1}$ and $v_{2}$ during round $t$ and $t-1$ while executing $\mathcal{A}$. So, at round $t$ since all whiteboard content are same for $R_{1}$ and $R_{2}$ (except for nodes $v_{1}$ and $v_{2}$ whose content only differs after round $t-2$ for $R_{1}$ and $R_{2}$ which can not be known by $a_{2}$ as it was not there during that rounds), they can not help $a_{2}$ to distinguish between $R_{1}$ and $R_{2}$. So the problem now can be thought of as solving PerpExploration-BBH with one agent where number of possible consecutive black hole position is at least 2. And this is impossible. So there doesn’t exist any algorithm that solves PerpExploration-BBH with two agents on a Ring $R$ where number of possible consecutive positions of black hole is at least 3.
Note that since the $n$ can be sufficiently large so that initially number of possible consecutive black hole positions is atleast three we have the following corollary.
@@
It is impossible to solve PerpExploration-BBH on a ring $R$ with $n$ nodes using two synchronous agents even if the nodes are equipped with whiteboards.
@@
A set of three synchronous agents are required to solve the PerpExploration-BBH on a ring $R$ with $n$ nodes, where each node has a whiteboard.
@@
A set of two agents, each equipped with $O(\log n)$ pebbles can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
Now the following theorem improves the lower bound for the case when agents are scattered and each agent is equipped with a pebble, whereas there is no whiteboard present at the nodes of $R$. The theorem is as follows.
@@
A set of 3 scattered agents, each equipped with a pebble can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
Proof 3.7.
Let $a_{1},a_{2}$ and $a_{3}$ be three agents equipped with a pebble each. The agents are placed on three nodes $h_{1},h_{2}$ and $h_{3}$ initially in such a way that distance between $h_{i}$ and $h_{j}$ is same for all $i,j\in\{1,2,3\}$, where we consider $h_{j}$ to be the nearest node of $h_{i}$ in the clockwise direction. Without loss of generality let this distances be sufficiently large. Further, let there exists an algorithm $\mathcal{A}$ that solves the PerpExploration-BBH problem in this setting. Let without loss of generality $a_{1}$ be the first agent to explore the third node from its corresponding starting position (i.e., $h_{1}$) in any of the clockwise or counter-clockwise direction, when each agent starts executing the algorithm $\mathcal{A}$. Suppose by following $\mathcal{A}$, $a_{1}$ can visit the third node in clockwise direction (without loss of generality) first at a round say $t>0$. Let $v_{1},v_{2}$ and $v_{3}$ be those set of three nodes from $h_{1}$ in the clockwise direction. Let $C_{1}$,$C_{2}$ and $C_{3}$ be three scenarios where in $C_{i}$, $v_{i}$ is the black hole. We claim that, $a_{1}$ can not carry its pebble during any execution of $\mathcal{A}$. Otherwise in scenario $C_{1}$, it would be consumed along with its pebble, and since the distances between two consecutive $h_{i}$ are sufficiently large, hence other agents would have no idea that an agent is already consumed. This is equivalent to solving the PerpExploration-BBH with two agents having a pebble each and as $n$ is sufficiently large this is impossible due to Corollary 3.5. Now suppose adversary chooses to activate the black hole whenever $a_{1}$ reaches there for the first time. In this case for all $C_{1}$, $C_{2}$ and $C_{3}$, at round $t$, agents $a_{2}$ and $a_{3}$ has no idea about where $a_{1}$ is consumed even if it knows that $a_{1}$ is consumed, as the distances between two consecutive $h_{i}$ are sufficiently large so their exploration region does not intersects till round $t$ and timeout (or,waiting for other agent) strategy do not work as for all $C_{i}$ an agent can get same timeout output. Also for all $C_{i}$s the position of the pebbles and alive agents will be same at round $t$. Now for the alive agents, the number of nodes for possible position of the black hole must be greater than or equal to 3 at round $t$. So the situation at round $t$ is similar to the problem of solving PerpExploration-BBH on a ring $R$ with two agents having a total of 3 pebbles where the number of possible consecutive positions of black hole is greater or equal to 3. Since we have assumed $\mathcal{A}$ solves the problem thus, $\mathcal{A}$ can also solve the problem of PerpExploration-BBH with two agents where number of possible positions of black hole is greater or equal to 3. But due to Theorem 3.1 it is impossible. Hence there can never exists any algorithm that solves PerpExploration-BBH with three scattered agents each of which are equipped with a pebble.
Corollary 3.8.
A set of four scattered agents each with a pebble, are required to solve PerpExploration-BBH on a ring $R$ with $n$ nodes.
4 Co-located Agents
4.1 Communication: Face-to-Face
In this section, we discuss the PerpExploration-BBH problem, under the conditions, where the agents are initially located at a safe node $home$, and they can only communicate with another agent whenever they are at the same node at the same round. Here considering this scenario we provide an algorithm PerpExplore-F2F (Algorithm 1) to solve the PerpExploration-BBH problem with $\lceil\log(n-1)\rceil+3$ agents.

### v2

@@ 3 Impossibility Results
Here, at beginning, we state the first impossibility result which gives us a lower bound on minimum number of agents required to solve PerpExploration-BBH.
@@
A set of two synchronous agents in a ring $R$ of size $n$ cannot solve PerpExploration-BBH, even in the presence of a whiteboard if number of possible consecutive black hole positions is at least 3.
The main idea of the proof of the above Theorem 3.1 is that we create a scenario using adversarial techniques, where after one agent is destroyed by the Byzantine black hole $v_{b}$ (say), the other and only alive agent ends up with at least two possible choices for the position of the black hole, which in turn creates confusion for the agent. Hence it is unable to correctly detect the black hole node among these two sets of nodes. In this scenario it is impossible for the only alive agent to successfully explore the whole ring except $v_{b}$ perpetually, either without correctly detecting the Byzantine black hole position, or without getting destroyed as well. As a direct implication of Theorem 3.1, we can have the following corollaries. (Please see Appendix in Section 7 for the detailed proof)
@@
A set of three synchronous agents are required to solve the PerpExploration-BBH on a ring $R$ with $n$ nodes, where each node of $R$ has a whiteboard.
Note that this lower bound of 3 agents is also true for agents with the pebble model of communication, as the pebble model of communication can be easily simulated in the whiteboard communication model.
@@
A set of two agents, each equipped with $O(\log n)$ pebbles, can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
Our next result further improves the lower bound for number of agents, when agents are scattered and have pebble model of communication.
Theorem 3.4.
A set of 3 scattered agents, each equipped with a pebble, can not solve the PerpExploration-BBH problem on a ring $R$ with $n$ nodes.
We have proved Theorem 3.4 using an adversarial technique. We created an instance of three copies of the same ring $R$ (i.e., $R_{1}$, $R_{2}$ and $R_{3}$) in such a way that the position of Byzantine black hole ($v_{b})$ are different in each of them but they are consecutive in $R$. Also, in each case, the agents are initially placed at the nodes of $R$, which are at a distance of equal length between each other. Now, when an agent is destroyed by $v_{b}$, the other two agents can not distinguish between these three copies (due to same information on nodes for each of them). So, the problem reduces to two agents with 3 pebbles solving PerpExploration-BBH where a number of possible consecutive black hole positions is at least three. Then by Corollary 3.3, we can have the desired result (For detailed proof, see Appendix in Section 8). Next, corollary is a direct consequence of Theorem 3.4.
@@
A set of four scattered agents, each with a pebble, are required to solve PerpExploration-BBH on a ring $R$ with $n$ nodes.
4 Perpetual Exploration with Co-located Agents
In this section we assume that initially agents are co-located at a node which is termed as home. On the basis of this assumption here we investigate the sufficiency for the number of agents to solve PerpExploration-BBH problem under different model of communication.
4.1 Pebble Model of Communication
In this section, we consider the communication model, where the starting node has $k-1$ identical and movable tokens (termed as pebbles), where $k$ is the total number of agents deployed. A pebble can be carried by an agent from one node to another. This pebble acts as a mode of communication for the agents, as the agents can perceive the presence of a pebble at the current node. Moreover, an agent can also perceive the presence of other agents which are co-located at the current round (i.e., gather the IDs of the other co-located agents). Our main aim is to prove the following theorem in this section.
@@ 4.1 Pebble Model of Communication
A team of 3 co-located, synchronous agents are necessary and sufficient to solve PerpExploration-BBH on a ring $R$, with $n$ nodes under the pebble model of communication,with the presence of two pebbles initially co-located with the agents.
The necessary part follows from Corollary 3.3. For the sufficiency part we present an algorithm, PerpExplore-Coloc-Pbl. We have provided a very brief description of the algorithm. For detailed description and pseudo code see Appendix (Section 9).